OpenStudy (anonymous):

Lets say we can drill a tunnel straight through north- and southpole. Describe the movement of a rock that falls from northpole. Lets say the earth has a homogenous density and discard friction. Calculate after how much time the stone reaches the mid-point of the earth.

OpenStudy (anonymous):

i know i should use : K = GmMr/R^3

OpenStudy (anonymous):

radius of the earth: 6371 km, mass: 5,9737.10^24

OpenStudy (festinger):

Just a suggestion: \[v=\frac{dy}{dt}\] Find an expression that defines the kinetic energy and thus speed at every point, then do an integral for y=R to 0. You should get an expression for time.

OpenStudy (anonymous):

find an expression the defines kinetic energy..

OpenStudy (anonymous):

you just mean the formula for kinet. en.?

OpenStudy (festinger):

yes.

OpenStudy (anonymous):

why? i have the answer of one of my classmates and it doesn't look like he used kinetic energy formula anywhere

OpenStudy (anonymous):

i see: \[\frac{ dm/dr*r }{ R^2 }\] and \[-kx=ma\]

OpenStudy (anonymous):

and the first equation i posted

OpenStudy (festinger):

I have no idea what \[\frac{ dm/dr*r }{ R^2 }\] is supposed to do, but if you want there is another way using -kx=ma. Recall what g is in terms of G, M and R. \[g=\frac{MG_{E}}{R^{2}}\] Thus, Force becomes: \[F=-mg\frac{r}{R}\] This looks similar to hooke's law:\[F=-kx\] Which you will find that:\[k=\frac{mg}{R}\] For spring oscillations,angular frequency is: \[\omega=\sqrt{\frac{k}{m}}\]\[T=2\pi\sqrt{\frac{mR}{mg}}\] This expression is for 1 compelte oscillation, so half of it is:\[t=\pi\sqrt{\frac{R}{g}}\]

OpenStudy (anonymous):

wow, alright

OpenStudy (fifciol):

I would do it like this: \[dy=\frac{ adt^2 }{ 2 }=\frac{ GMdt^2 }{ 2y^2 }\] \[2dyy^2=GMdt^2 \rightarrow 2 \int\limits_{0}^{R}y^2dy=GM \int\limits_{0}^{t}dt^2\] \[\frac{ 2R^3 }{ 3 }=GMt^2 \rightarrow t=\sqrt{\frac{ 2R^3 }{3GM }}=657,65 s\]

OpenStudy (anonymous):

could you explain hooke's law? what is k and x?

OpenStudy (festinger):

Well, you threw out the idea of :−kx=ma... Hooke's law states that the restoring force is proportional to the displacement x. this proportionality constant is k. @Fifciol i'll try to make sense of that :P

OpenStudy (anonymous):

restoring force..proportionality constant... alright...

OpenStudy (anonymous):

@Fifciol you used the kinematic equation for dy?

OpenStudy (anonymous):

then dt^2 so it's a function of time...?

OpenStudy (fifciol):

I used kinematic equation and dt^2 is differential

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

and a = GM/y^2 ???

OpenStudy (fifciol):

yes

OpenStudy (anonymous):

how'd you think of that?

OpenStudy (festinger):

F=ma=mg

OpenStudy (fifciol):

F=GMm/y^2=ma

OpenStudy (anonymous):

y is the radius, right?

OpenStudy (fifciol):

distance

OpenStudy (fifciol):

R would be at surface of the earth

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

so you described the movement using kinematic equations, then integrated to know the time?

OpenStudy (fifciol):

yes

OpenStudy (anonymous):

got it, i think it's correct. Unless @Festinger has a different answer?

OpenStudy (fifciol):

I'm not sure the first equation I wrote down

OpenStudy (festinger):

Even my first suggestion gave me a different expression.

OpenStudy (anonymous):

@Mashy any thoughts?

OpenStudy (fifciol):

@amistre64 is that correct: \[dy=\frac{ adt^2 }{ 2 }\] ?

OpenStudy (anonymous):

if a is constant, it is right?

OpenStudy (amistre64):

i havent really learned to model differential equations

OpenStudy (amistre64):

wouldnt this act like a pendulum tho? or are we trying to determine when itll first reach the center?

OpenStudy (anonymous):

when it will reach the center

OpenStudy (festinger):

I am unsure if \[\int_{0}^{t}dt^{2}=t^{2}\] is a legal move....

OpenStudy (amistre64):

and there is some relation to the inverse of a square of a distance right?

OpenStudy (anonymous):

Fificiol is wrong.. cause the mass of the earth is not a constant as you go down the earth..

OpenStudy (anonymous):

i mean the effective mass that provides the gravitational force :P

OpenStudy (amistre64):

doesnt assuming homogenous density and whatnot mean that we are simplifying all the complexity from it?

OpenStudy (anonymous):

yesyes

OpenStudy (anonymous):

yes.. so u need to get mass as a function of distance.. use the densitiy as a constant.. and get a function for mass in terms of M and R

OpenStudy (fifciol):

it has homogenous density, and when you deal with gravity and planets you can treat the mass as if all of it were at center

OpenStudy (festinger):

While true that:\[\int d(t^{2})=t^{2}+c\] is true, but in this case it is actually: \[\int (dt)^{2}\]

OpenStudy (anonymous):

@fificiol.. as you go inside the earth, the gravitational force decreases.. in your equation u should modify M m = My^3/R^3...

OpenStudy (anonymous):

because according to your equation, the force keeps increasing as you go inside the earth.. so you get a huge error.. u ll get acceleration to be increasing.. infact the acceleration is decreasing! ^_^

OpenStudy (anonymous):

i think.. everything else holds..!!

OpenStudy (festinger):

Mashy is right.

OpenStudy (anonymous):

i'm seeing this equation from my classmate: \[m.a=\frac{ G.m.M }{ R^3 }\]

OpenStudy (anonymous):

times r

OpenStudy (fifciol):

yes, Mashy you're right

OpenStudy (anonymous):

\[m.a=\frac{ G.m.M }{ R^3 } . r\]

OpenStudy (anonymous):

so...what now?

OpenStudy (festinger):

follow what i posted for now until someone comes with a better solution. i am uncomfortable with the pi appearing...

OpenStudy (anonymous):

ok

OpenStudy (festinger):

but i would really suggest you go back to look at hooke's law. your friend seems to know hooke's law.

OpenStudy (anonymous):

i don't understand how it applies to this

OpenStudy (festinger):

the expression that you posted that your friend wrote ma=-kx gave me the idea to link hooke's law and gravity in this case. the equation of force shows that it is proportional to r, like the x that appears in hooke's law. after that i just rode on the results from hooke's law with spring oscillations.

OpenStudy (anonymous):

i see, you used it for the proportionality

OpenStudy (fifciol):

My new way of thinking on this problem:  \[ma=-kx \rightarrow x''+\frac{ k }{m }x=0\] \[\frac{ GMm }{ R^2 }=kR \rightarrow k=\frac{ GMm }{ R^3 }\] \[x''+\frac{ GM }{ R^3 }x=0 \rightarrow T=2\pi \sqrt{\frac{ R^3 }{ GM }}\] \[t=\frac{T }{ 4 }=\frac{ \pi }{ 2 }\sqrt{\frac{ R^3 }{ GM }}\]

OpenStudy (festinger):

Yes, you are right. It should be divided by 4 instead of 2. My bad.

OpenStudy (anonymous):

why?

OpenStudy (festinger):

One oscillation is from one end > center > another end > center > original end when i divided by 2 it is the solution to: one end > center > another end

OpenStudy (anonymous):

and you need, one end>center ?

OpenStudy (festinger):